### v1

# Decoupled structure-preserving discretization of incompressible MHD equations with general boundary conditions

Abstract
In the framework of a mixed finite element method, a structure-preserving formulation for incompressible MHD equations with general boundary conditions is proposed. A leapfrog-type temporal scheme fully decouples the fluid part from the Maxwell part by means of staggered discrete time sequences and, in doing so, partially linearizes the system. Conservation and dissipation properties of the formulation before and after the decoupling are analyzed. We demonstrate optimal spatial and second-order temporal error convergence and conservation and dissipation properties of the proposed method using manufactured solutions, and apply it to the benchmark Orszag-Tang and lid-driven cavity test cases.
keywords:
MHD , mixed finite element method , structure-preserving discretization , general boundary conditions
journal: arXiv
[…]
[1]organization=School of Mathematics and Computing Science, Guilin University of Electronic Technology, city=Guilin, country=China
[…]
[2]organization=Delft Institute of Applied Mathematics, Delft University of Technology, city=Delft, country=the Netherlands
[…]
[3]organization=ICA, Université de Toulouse, ISAE–SUPAERO, INSA, CNRS, MINES ALBI, UPS, city=Toulouse, country=France
[…]
[4]organization=Faculty of Aerospace Engineering, Delft University of Technology, city=Delft, country=the Netherlands
1 Introduction
Given a bounded, contractible domain $\Omega\subset\mathbb{R}^{d}$, $d\in\left\{2,3\right\}$ with a Lipschitz boundary $\partial\Omega$, the incompressible, constant density magnetohydrodynamic (or simply incompressible MHD) equations [1] written as
(1a) $\rho\left[\partial_{t}\bm{u}+\left(\bm{u}\cdot\nabla\right)\bm{u}\right]-\tilde{\mu}\Delta\bm{u}-\bm{j}\times\bm{B}+\nabla p$ $=\rho\bm{f},$
(1b) $\nabla\cdot\bm{u}$ $=0,$
(1c) $\partial_{t}\bm{B}+\nabla\times\bm{E}$ $=\bm{0},$
(1d) $\bm{j}-\sigma\left(\bm{E}+\bm{u}\times\bm{B}\right)$ $=\bm{0},$
(1e) $\bm{j}-\nabla\times\bm{H}$ $=\bm{0},$
(1f) $\bm{B}$ $=\mu\bm{H},$
govern the dynamics of fluid velocity $\bm{u}$, electric current density $\bm{j}$, magnetic flux density $\bm{B}$, magnetic field strength $\bm{H}$, hydrodynamic pressure $p$ and electric field strength $\bm{E}$, subject to body force field $\bm{f}$, a velocity initial condition $\bm{u}^{0}:=\bm{u}(\bm{x},0)$ and a divergence-free magnetic initial condition $\bm{B}^{0}:=\bm{B}(\bm{x},0)$, and proper boundary conditions [2], in a space-time domain, $\Omega\times(0,T]$. The material parameters involved are the fluid density $\rho$, the dynamic viscosity $\tilde{\mu}$, the electric conductivity $\sigma$, and the magnetic permeability $\mu$.
The incompressible MHD equations, (1), are a coupled system of the incompressible Navier-Stokes equations and the Maxwell’s equations. The two evolution equations, (1a) and (1c), stand for the conservation of momentum and the Faraday’s law of induction, respectively. Ohm’s law and Ampere’s law are (1d) and (1e), respectively. (1f) sets up a constitutive relation between $\bm{B}$ and $\bm{H}$. The divergence-free constraint on $\bm{u}$, (1b), represents mass conservation of the constant mass density flow. It is seen that, as $\bm{B}^{0}:=\bm{B}(\bm{x},0)$ is divergence free, (1c) ensures
$\nabla\cdot\bm{B}=0\quad\mathrm{in}\ \Omega\times(0,T],$
which is Gauss’s law for magnetism. And, from (1e), we know $\nabla\cdot\bm{j}=0$ which stands for conservation of charge. The coupling between the flow field and the electromagnetic field is through the Lorentz force, $\bm{j}\times\bm{B}$ in (1a), and the electromotive force, $\bm{u}\times\bm{B}$ in (1d).
Let $L$, $U$, and $B$ denote characteristic quantities of length, velocity, and magnetic flux density, respectively. We scale the space-time domain and variables by
$\bm{x}\leftarrow\dfrac{\bm{x}}{L},\quad t\leftarrow\dfrac{Ut}{L},$
$\bm{u}\leftarrow\dfrac{\bm{u}}{U},\quad p\leftarrow\dfrac{p}{\rho U^{2}},\quad\bm{f}\leftarrow\dfrac{L\bm{f}}{U^{2}},$
$\bm{B}\leftarrow\dfrac{\bm{B}}{B},\quad\bm{H}\leftarrow\dfrac{\mu\bm{H}}{B},\quad\bm{E}\leftarrow\dfrac{\bm{E}}{UB},\quad\bm{j}\leftarrow\dfrac{\mu L\bm{j}}{B}.$
And the induced scaled operators are
$D\leftarrow LD,\ D\in\left\{\nabla,\nabla\times,\nabla\cdot\right\},$
$\Delta\leftarrow L^{2}\Delta,\quad\partial_{t}\leftarrow\dfrac{L}{U}\partial_{t},\quad\bm{u}\cdot\nabla\leftarrow\dfrac{L}{U}\left(\bm{u}\cdot\nabla\right).$
Note that we do not introduce new notations for the scaled objects and, after this point, they always refer to the scaled dimensionless ones.
Now (1) can be written in a dimensionless form, i.e.,
(2a) $\partial_{t}\bm{u}+\bm{\omega}\times\bm{u}+\mathrm{R}_{f}^{-1}\nabla\times\bm{\omega}-\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}+\nabla P$ $=\bm{f},$ in $\qquad\text{in }\Omega\times(0,T],$
(2b) $\bm{\omega}-\nabla\times\bm{u}$ $=\bm{0},$ in $\qquad\text{in }\Omega\times(0,T],$
(2c) $\nabla\cdot\bm{u}$ $=0,$ in $\qquad\text{in }\Omega\times(0,T],$
(2d) $\partial_{t}\bm{H}+\mathrm{R}_{m}^{-1}\Delta\bm{H}-\nabla\times\left(\bm{u}\times\bm{H}\right)$ $=\bm{0},$ in $\qquad\text{in }\Omega\times(0,T],$
[…]
(3a) $\partial_{t}\bm{B}+\nabla\times\bm{E}$ $=\bm{0},$
(3b) $\mathrm{R}_{m}^{-1}\bm{j}-\left(\bm{E}+\bm{u}\times\bm{B}\right)$ $=\bm{0},$
(3c) $\bm{j}-\nabla\times\bm{H}$ $=\bm{0},$
(3d) $\bm{B}$ $=\bm{H},$
have been written into (2d), which is the magnetic induction (or Helmholtz) equation. [^1] Note that, after non-dimensionalization, we get $\bm{B}=\bm{H}$, i.e. (3d). This is the case in vector calculus. In differential forms, they appear as a pair of differently oriented dual forms connected by the Hodge $\star$ operator, i.e., $\bm{B}=\star\bm{H}$. The three dimensionless parameters are the fluid Reynolds number $\mathrm{R}_{f}=\dfrac{\rho UL}{\tilde{\mu}}=\dfrac{UL}{\nu}$ (with $\nu=\dfrac{\tilde{\mu}}{\rho}$ being the kinematic viscosity), the magnetic Reynolds number $\mathrm{R}_{m}=\mu\sigma UL$, and the coupling number $\mathsf{c}=\mathrm{A}_{l}^{-2}$ ($\mathrm{A}_{l}=\dfrac{U\sqrt{\rho\mu}}{B}=\dfrac{U}{U_{\mathrm{A}}}$ is the Alfvén number and $U_{\mathrm{A}}=\dfrac{B}{\sqrt{\rho\mu}}$ is the Alfvén speed). Sometimes, the coupling term in (2a), $\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}$, is written as $\mathsf{s}\left(\bm{E}+\bm{u}\times\bm{H}\right)\times\bm{H}$ where the coupling number becomes $\mathsf{s}=\mathsf{c}\mathrm{R}_{m}=\dfrac{\sigma LB^{2}}{\rho U}$. For example, see [3, (2.2a)].
To close the problem, (2) is complemented with initial conditions $\bm{u}^{0},\bm{H}^{0}$ and boundary conditions. In this work, we consider the following general boundary condition configuration:
  • 1.
    For the fluid part, there are two pairs of boundary conditions,
    (4) {$&P=\widehat{P}\quad&&\text{on }\Gamma_{\widehat{P}}\times\left(0,T\right]\\ &\bm{u}\cdot\bm{n}=\widehat{u}\quad&&\text{on }\Gamma_{\widehat{u}}\times\left(0,T\right]\end{aligned}\right.\quad\text{and}\quad\left\{\begin{aligned} &\bm{u}\times\bm{n}=\widehat{\bm{u}}\quad&&\text{on }\Gamma_{\widehat{\bm{u}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{\omega}\times\bm{n}\right)=\widehat{\bm{\omega}}\quad&&\text{on }\Gamma_{\widehat{\bm{\omega}}}\times\left(0,T\right]\end{aligned}\right.,$
    where $\bm{n}$ is the outward unit normal vector, $\left\{\Gamma_{\widehat{P}},\Gamma_{\widehat{u}}\right\}$ and $\left\{\Gamma_{\widehat{\bm{u}}},\Gamma_{\widehat{\bm{\omega}}}\right\}$ are two partitions of the boundary, such that $\overline{\Gamma}_{\widehat{P}}\cup\overline{\Gamma}_{\widehat{u}}=\partial\Omega$, $\Gamma_{\widehat{P}}\cap\Gamma_{\widehat{u}}=\emptyset$ and $\overline{\Gamma}_{\widehat{\bm{u}}}\cup\overline{\Gamma}_{\widehat{\bm{\omega}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{u}}}\cap\Gamma_{\widehat{\bm{\omega}}}=\emptyset$.
  • 2.
    For the Maxwell part, boundary conditions are given as
    (5) $&\bm{E}\times\bm{n}=\widehat{\bm{E}}\quad&&\text{on }\Gamma_{\widehat{\bm{E}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{H}\times\bm{n}\right)=\widehat{\bm{H}}\quad&&\text{on }\Gamma_{\widehat{\bm{H}}}\times\left(0,T\right]\end{aligned}\right.,$
    where $\left\{\Gamma_{\widehat{\bm{E}}},\Gamma_{\widehat{\bm{H}}}\right\}$ is a partition of the boundary, satisfying $\overline{\Gamma}_{\widehat{\bm{E}}}\cup\overline{\Gamma}_{\widehat{\bm{H}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{E}}}\cap\Gamma_{\widehat{\bm{H}}}=\emptyset$. Note that the boundary condition $\bm{E}\times\bm{n}$ implies the magnetic flux boundary condition $\bm{H}\cdot\bm{n}$ (or $\bm{B}\cdot\bm{n}$). This is because
[…]
    on $\partial\Omega\times\left(0,T\right]$. Also see [4, Section 2.3].
For example, the boundary conditions considered in the work of Hu et al., see (7) of [5], are a particular case of this general configuration, i.e. the case when $\Gamma_{\widehat{u}}=\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset$, $\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$.
It is generally acknowledged that structure-preservation, especially the preservation (either strongly of weakly) of $\nabla\cdot\bm{B}=0$, is an essential property of numerical methods for MHD [6]. Many of these structure-preserving methods can be classified into, for example, potential-based methods [7], divergence-cleaning methods [8, 9], constrained transport methods [10, 11] and Lagrange multiplier (or augmented) methods [12, 13, 14]. In the current paper neither of those techniques will be applied, but it will be shown that the absence of magnetic monopoles is weakly enforced through the decoupled variational formulation. Recently, methods that aim at preserving multiple structures in the mixed finite element setting become increasingly popular. See, for example, the work of Hu et al. for incompressible MHD that preserves cross- and magnetic-helicity, energy and Gauss law of magnetism [5], the work of Gawlik and Gay-Balmaz for incompressible MHD of a variable fluid density that preserves energy, cross-helicity (when the fluid density is constant) and magnetic-helicity, mass, total squared density, pointwise incompressibility, and Gauss’s law for magnetism [15], the work of Laakmann et al. [16] and references therein. These methods preserve several physical quantities of interest, but usually are less computationally efficient because of their large (nonlinear) discrete systems to be solved. A compromise is to use a decoupled scheme which breaks the discrete system into several smaller ones, for example see [3, 12, 13, 17, 18]. These methods in the mixed finite element setting commonly deal with a particular configuration of boundary conditions.
In this paper, we introduce a structure-preserving mixed finite element discretization for the incompressible MHD under a general boundary condition configuration. The method is decoupled in the sense that the two evolution equations are temporally discretized at different time sequences such that they can be solved individually.
The outline of the rest of the paper is as follows. Relevant Hilbert spaces and the de Rham Hilbert complex are introduced in Section 2 followed by the proposed semi-discrete formulation and its conservation properties in Section 3. The temporal discretization that leads to the decoupled scheme is explained in Section 4. Numerical tests are presented in Section 5. Finally, in Section 6 conclusions are drawn.
2 Relevant Hilbert spaces and the de Rham complex
The function spaces used in this work are similar to those used in [19] where only the two-dimensional setting is discussed. For the sake of completeness, we give a brief introduction to the three-dimensional spaces as a complement to [19, Section 2.1]. We first introduce the relevant spaces and the de Rham complex in the infinite dimensional setting, and then introduce their finite dimensional counterparts. For an extensive introduction of function spaces, see [20] or any other textbook on functional analysis.
2.1 The infinite dimensional setting
Let $L^{2}\left(\Omega\right)$ denote the space of square-integrable functions on $\Omega$,
$L^{2}(\Omega):=\left\{\varphi\left|\left\langle{\varphi},{\varphi}\right\rangle_{\Omega}:=\int_{\Omega}\varphi\cdot\varphi\ \mathrm{d}\Omega<\infty\right.\right\}.$
For $\Omega\subset\mathbb{R}^{3}$, we also consider the following Hilbert spaces,
$H^{1}\left(\Omega\right)$ :=$:=\left\{\psi\left|\psi\in L^{2}(\Omega),\nabla\psi\in\left[L^{2}\left(\Omega\right)\right]^{3}\right.\right\},$
$H\left(\mathrm{curl};\Omega\right)$ :=$:=\left\{\bm{\sigma}\left|\bm{\sigma}\in\left[L^{2}\left(\Omega\right)\right]^{3},\nabla\times\bm{\sigma}\in\left[L^{2}\left(\Omega\right)\right]^{3}\right.\right\},$
$H\left(\mathrm{div};\Omega\right)$ :=$:=\left\{\bm{\phi}\left|\bm{\phi}\in\left[L^{2}\left(\Omega\right)\right]^{3},\nabla\cdot\bm{\phi}\in L^{2}\left(\Omega\right)\right.\right\}.$
They form the well known de Rham Hilbert complex [21, 22],
(7) $\mathbb{R}\hookrightarrow H^{1}\left(\Omega\right)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}H(\mathrm{curl};\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}H(\mathrm{div};\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}L^{2}(\Omega)\to 0.$
[…]
The trace operator, $\mathcal{T}$, restricts an element of Hilbert spaces to a boundary section $\Gamma\subseteq\partial\Omega$. For $\psi\in H^{1}\left(\Omega\right)$ and $\bm{\phi}\in H\left(\mathrm{div};\Omega\right)$, we define its action as
$\mathcal{T}\psi=\left.\psi\right|_{\Gamma},\quad\mathcal{T}\bm{\phi}=\left.\bm{\phi}\cdot\bm{n}\right|_{\Gamma}.$
And we distinguish between $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ for elements of $H\left(\mathrm{curl};\Omega\right)$, i.e., for $\bm{\sigma},\bm{\eta}\in H\left(\mathrm{curl};\Omega\right)$,
(8) $\mathcal{T}\bm{\sigma}=\left.\bm{\sigma}\times\bm{n}\right|_{\Gamma},\quad\mathcal{T}_{\parallel}\bm{\eta}=\left.\bm{n}\times\left(\bm{\eta}\times\bm{n}\right)\right|_{\Gamma}.$
Using the above, we define the trace spaces on $\Gamma$ as
$H^{1/2}\left(\Omega,\Gamma\right):=\left\{\mathcal{T}\psi\left|\psi\in H^{1}(\Omega)\right.\right\},$
$\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right):=\left\{\mathcal{T}\bm{\sigma}\left|\bm{\sigma}\in H(\mathrm{curl};\Omega)\right.\right\},$
$\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right):=\left\{\mathcal{T}_{\parallel}\bm{\eta}\left|\bm{\eta}\in H(\mathrm{curl};\Omega)\right.\right\},$
$\mathcal{T}H\left(\mathrm{div};\Omega,\Gamma\right):=\left\{\mathcal{T}\bm{\phi}\left|\bm{\phi}\in H(\mathrm{div};\Omega)\right.\right\}.$
Note that, at the continuous level, although $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ are different, see (8), trace spaces $\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right)$ and $\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right)$ refer to the same tangential space.
2.2 The finite dimensional setting
Assume that we are given finite dimensional spaces $G(\Omega)\subset H^{1}(\Omega)$, $C(\Omega)\subset H(\mathrm{curl};\Omega)$, $D(\Omega)\subset H(\mathrm{div};\Omega)$, $S(\Omega)\subset L^{2}(\Omega)$, which form a discretization of the continuous de Rham Hilbert complex,
(9) $\mathbb{R}\hookrightarrow G(\Omega)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}C(\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}D(\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}S(\Omega)\to 0.$
In particular, this means that if the complex in (7) is exact, then so is the one in (9). We will also consider the following subspaces with boundary conditions, where $\Gamma\subset\partial\Omega$,
$G_{\widehat{\psi}}(\Omega,\Gamma):=\left\{\psi_{h}\left|\psi_{h}\in G(\Omega),\ \mathcal{T}\psi_{h}=\widehat{\psi}\ \text{on}\ \Gamma\right.\right\},$
$C_{\bm{\widehat{\sigma}}}(\Omega,\Gamma):=\left\{\bm{\sigma}_{h}\left|\bm{\sigma}_{h}\in C(\Omega),\ \mathcal{T}\bm{\sigma}_{h}=\bm{\widehat{\sigma}}\ \text{on}\ \Gamma\right.\right\},$
$C^{\parallel}_{\bm{\widehat{\eta}}}(\Omega,\Gamma):=\left\{\bm{\eta}_{h}\left|\bm{\eta}_{h}\in C(\Omega),\ \mathcal{T}_{\parallel}\bm{\eta}_{h}=\bm{\widehat{\eta}}\ \text{on}\ \Gamma\right.\right\},$
$D_{\widehat{\phi}}(\Omega,\Gamma):=\left\{\bm{\phi}_{h}\left|\bm{\phi}_{h}\in D(\Omega),\ \mathcal{T}\bm{\phi}_{h}=\widehat{\phi}\ \text{on}\ \Gamma\right.\right\}.$
For the corresponding homogeneous boundary conditions we have
$G_{0}(\Omega,\Gamma):=\left\{\psi_{h}\left|\psi_{h}\in G(\Omega),\ \mathcal{T}\psi_{h}=0\ \text{on}\ \Gamma\right.\right\},$
$C_{\bm{0}}(\Omega,\Gamma):=\left\{\bm{\sigma}_{h}\left|\bm{\sigma}_{h}\in C(\Omega),\ \mathcal{T}\bm{\sigma}_{h}=\bm{0}\ \text{on}\ \Gamma\right.\right\},$
$C^{\parallel}_{\bm{0}}(\Omega,\Gamma):=\left\{\bm{\eta}_{h}\left|\bm{\eta}_{h}\in C(\Omega),\ \mathcal{T}_{\parallel}\bm{\eta}_{h}=\bm{0}\ \text{on}\ \Gamma\right.\right\},$
$D_{0}(\Omega,\Gamma):=\left\{\bm{\phi}_{h}\left|\bm{\phi}_{h}\in D(\Omega),\ \mathcal{T}\bm{\phi}_{h}=0\ \text{on}\ \Gamma\right.\right\}.$
Moreover, the finite dimensional trace spaces will be denoted by $\mathcal{T}G(\Omega,\Gamma)$, $\mathcal{T}C(\Omega,\Gamma)$, $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$, $\mathcal{T}D(\Omega,\Gamma)$. Unlike their continuous counterparts, finite dimensional trace spaces $\mathcal{T}C(\Omega,\Gamma)$ and $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$ usually are different.
We define a corresponding trilinear form,
(10) $a\left(\bm{\alpha}_{h},\bm{\beta}_{h},\bm{\gamma}_{h}\right):=\left\langle{\bm{\alpha}_{h}\times\bm{\beta}_{h}},{\bm{\gamma}_{h}}\right\rangle_{\Omega},\quad\bm{\alpha}_{h},\bm{\beta}_{h},\bm{\gamma}_{h}\in\left\{C(\Omega),D(\Omega)\right\},$
which is skew-symmetric with respect to any two of the three entries. Note that, in the infinite dimensional setting, the $L^{2}$-boundedness of $\left\langle{\bm{\alpha}\times\bm{\beta}},{\bm{\gamma}}\right\rangle_{\Omega}$ is not implied even if $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are $L^{2}$ vector fields. However, in the finite dimensional setting where the spaces $C(\Omega)$ and $D(\Omega)$ consist of piecewise-polynomial finite element vector fields, it is in fact $L^{2}$-bounded.
3 Spatially-discrete formulation and its conservation properties
In this section, we present a spatially-discrete weak formulation of (2) and analyze its conservation properties. As a remark, for neatness, we only discuss the three-dimensional formulation and its temporal discretization in this paper. Obtaining their two-dimensional versions is straightforward.
3.1 Spatially-discrete formulation
To increase the notational clarity, we omit the part for the temporal domain. The proposed semi-discrete weak formulation of (2) in $\mathbb{R}^{3}$ is written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, seek $\left(\bm{u}_{h},\bm{\omega}_{h},P_{h},\bm{H}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}_{\perp}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
(11a) $\left\langle\partial_{t}\bm{u}_{h},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\bm{\omega}_{h},\bm{u}_{h},\bm{v}_{h}\right)+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\bm{\omega}_{h}},{\bm{v}_{h}}\right\rangle_{\Omega}\qquad$
$-\mathsf{c}\ a\left(\nabla\times\bm{H}_{h},\bm{H}_{h},\bm{v}_{h}\right)-\left\langle{P_{h}},{\nabla\cdot\bm{v}_{h}}\right\rangle_{\Omega}$ $=\left\langle{\bm{f}},{\bm{v}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}},{\mathcal{T}\bm{v}_{h}}\right\rangle_{\Gamma_{\widehat{P}}},$
(11b) $-\left\langle{\bm{u}_{h}},{\nabla\times\bm{w}_{h}}\right\rangle_{\Omega}+\left\langle{\bm{\omega}_{h}},{\bm{w}_{h}}\right\rangle_{\Omega}$ $=-\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{w}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{u}}}},$
(11c) $\left\langle{\nabla\cdot\bm{u}_{h}},{q_{h}}\right\rangle_{\Omega}$ $=0,$
(11d) $\left\langle{\partial_{t}\bm{H}_{h}},{\bm{b}_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\bm{b}_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\bm{b}_{h}\right)$ $=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{b}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}.$
Note that this assumes that the supplied boundary conditions are exactly representable as members of the chosen finite dimensional spaces; more generally, one can always project the boundary conditions into the appropriate finite dimensional trace spaces. See (4) and (5) for the general boundary setting.
3.2 Conservation properties
In this section, we analyze some properties of the formulation (11). In particular, we show that this semi-discrete formulation preserves conservation of mass and conservation of charge strongly, preserves Gauss’ law of magnetism weakly, and preserves the correct energy dissipation rate such that, when no body forces are applied and there is no net flux of energy through the boundary, it conserves the total energy in the ideal limit.
3.2.1 Conservation of mass
[…]
3.2.2 Conservation of charge
Since we have selected $\bm{H}_{h}\in C(\Omega)$, from (9), we know that we can always find $\bm{j}_{h}=\nabla\times\bm{H}_{h}\in D(\Omega)$ which is pointwise divergence-free; conservation of charge is preserved by the formulation (11).
3.2.3 Weak conservation of Gauss’s law for magnetism
Since (11d) is valid for all $\bm{b}_{h}\in C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$, let us choose $\bm{b}_{h}=\nabla q_{h}$ for an arbitrary $q_{h}\in G_{0}(\Omega,\Gamma_{\widehat{\bm{H}}})$. Then, we have
(12) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\nabla q_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\nabla q_{h}\right)=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\nabla q_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\;.$
The second and third terms on the left hand side are zero because we strongly satisfy ×$\nabla\times\nabla q_{h}=0$ since our discrete function spaces constitute a finite dimensional de Rham complex. We are left with the following equation,
(13) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\nabla q_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\,.$
Using the fact that $\nabla q_{h}\in C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$, we have
(14) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}=\left\langle\widehat{\bm{E}},\mathcal{T}_{\parallel}\nabla q_{h}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}$ $=\left\langle\widehat{\bm{E}},\mathcal{T}_{\parallel}\nabla q_{h}\right\rangle_{\partial\Omega}$
$=-\left\langle\left(\nabla\cdot\right)_{\partial\Omega}\widehat{\bm{E}},\mathcal{T}q_{h}\right\rangle_{\partial\Omega}=-\left\langle\left(\nabla\cdot\right)_{\Gamma_{\widehat{\bm{E}}}}\widehat{\bm{E}},\mathcal{T}q_{h}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\;.$
Finally, using the relation between $\bm{H}\cdot\bm{n}$ and $\widehat{\bm{E}}$ from (6), we can rewrite the rightmost term to obtain the following equation,
(15) $\dfrac{\mathrm{d}}{\mathrm{d}t}\left[\left\langle{\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}-\left\langle\bm{H}\cdot\bm{n},\mathcal{T}q_{h}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\right]=0\;.$
The expression inside the square brackets is precisely the weak divergence-free constraint on $\bm{H}_{h}$. The above equation thus states that the time-derivative of this constraint on $\bm{H}_{h}$ is zero, and implies that $\bm{H}_{h}$ will remain weakly divergence-free for all $t$ if the initial condition $\bm{H}^{0}_{h}$ is weakly divergence-free.
3.2.4 Energy Conservation
The (semi-)discrete (total) energy is defined as
(16) $\mathcal{E}_{h}:=\mathcal{K}_{h}+\mathcal{M}_{h}$
whose $\mathcal{K}_{h}:=\dfrac{1}{2}\left\langle{\bm{u}_{h}},{\bm{u}_{h}}\right\rangle_{\Omega}$ is the kinetic energy of the fluid and $\mathcal{M}_{h}:=\dfrac{\mathsf{c}}{2}\left\langle{\bm{H}_{h}},{\bm{H}_{h}}\right\rangle_{\Omega}$ is the magnetic energy. Suppose
(17) $\widehat{u}=0\;,\quad\Gamma_{\widehat{\bm{u}}}=\Gamma_{\widehat{\bm{E}}}=\partial\Omega\;,\quad\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset\;.$
By replacing $\bm{v}_{h}$ in (11a) with $\bm{u}_{h}$, we can obtain
(18) $\partial_{t}\mathcal{K}_{h}+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\bm{\omega}_{h}},{\bm{u}_{h}}\right\rangle_{\Omega}-\mathsf{c}\ a\left(\bm{j}_{h},\bm{H}_{h},\bm{u}_{h}\right)=\left\langle{\bm{f}},{\bm{u}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}},{\mathcal{T}\bm{u}_{h}}\right\rangle_{\Gamma_{\widehat{P}}},$
where $\bm{j}_{h}=\nabla\times\bm{H}_{h}\in D(\Omega)$, and the fluid advection term $a\left(\bm{\omega}_{h},\bm{u}_{h},\bm{u}_{h}\right)$ and the total pressure term $\left\langle{P_{h}},{\nabla\cdot\bm{u}_{h}}\right\rangle_{\Omega}$ have vanished because of the skew-symmetry of the trilinear (10) and $\nabla\cdot\bm{u}_{h}=0$, respectively. Meanwhile, because (11b) holds $\forall\bm{w}_{h}\in C(\Omega)$ and $\bm{\omega}_{h}\in C(\Omega)$, we know that
$-\left\langle{\bm{u}_{h}},{\nabla\times\bm{\omega}_{h}}\right\rangle_{\Omega}+\left\langle{\bm{\omega}_{h}},{\bm{\omega}_{h}}\right\rangle_{\Omega}=-\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{\omega}_{h}}\right\rangle_{\partial\Omega}.$
As a results, (18) can be expressed as
(19) $\partial_{t}\mathcal{K}_{h}+\mathrm{R}_{f}^{-1}\left\langle{\bm{\omega}_{h}},{\bm{\omega}_{h}}\right\rangle_{\Omega}-\mathsf{c}\ a\left(\bm{j}_{h},\bm{H}_{h},\bm{u}_{h}\right)=\left\langle{\bm{f}},{\bm{u}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}},{\mathcal{T}\bm{u}_{h}}\right\rangle_{\Gamma_{\widehat{P}}}+\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{\omega}_{h}}\right\rangle_{\partial\Omega}.$
Similarly, replacing $\bm{b}_{h}$ in (11d) with $\bm{H}_{h}$ leads to
(20) $\dfrac{1}{\mathsf{c}}\partial_{t}\mathcal{M}_{h}+\mathrm{R}_{m}^{-1}\left\langle{\bm{j}_{h}},{\bm{j}_{h}}\right\rangle_{\Omega}-a\left(\bm{u}_{h},\bm{H}_{h},\bm{j}_{h}\right)=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{H}_{h}}\right\rangle_{\partial\Omega}.$
With (19) and (20), we can finally derive the rate of change of the discrete energy,
(21) $\partial_{t}\mathcal{E}_{h}$ $=\partial_{t}\mathcal{K}_{h}+\partial_{t}\mathcal{M}_{h}$
$=\underbrace{\left\langle{\bm{f}},{\bm{u}_{h}}\right\rangle_{\Omega}}_{\text{(i)}}\ \underbrace{-\mathrm{R}_{f}^{-1}\mathcal{S}_{h}}_{\text{(ii)}}\ \underbrace{-\mathsf{c}\mathrm{R}_{m}^{-1}\mathcal{J}_{h}}_{\text{(iii)}}\underbrace{-\left\langle{\widehat{P}},{\mathcal{T}\bm{u}_{h}}\right\rangle_{\Gamma_{\widehat{P}}}}_{\text{(iv)}}+\underbrace{\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{\omega}_{h}}\right\rangle_{\partial\Omega}}_{\text{(v)}}+\underbrace{\mathsf{c}\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{H}_{h}}\right\rangle_{\partial\Omega}}_{\text{(vi)}},$
where $\mathcal{S}_{h}:=\left\langle{\bm{\omega}_{h}},{\bm{\omega}_{h}}\right\rangle_{\Omega}$, $\mathcal{J}_{h}:=\left\langle{\bm{j}_{h}},{\bm{j}_{h}}\right\rangle_{\Omega}$. The two nonlinear terms have canceled each other due to the skew-symmetry of the trilinear form (10); the exchange of energy between $\mathcal{K}_{h}$ and $\mathcal{M}_{h}$ is exactly captured.
In (21), (i) represents amount of mechanical work done on the system per unit time, (ii) and (iii) are dissipation rates due to viscosity and electric resistance, respectively, and terms (iv) - (vi) are the net flux of energy through the domain boundary.
Clearly, if $\bm{f}=\bm{0}$, when the MHD flow is ideal (i.e. $\mathrm{R}_{f}=\mathrm{R}_{m}=\infty$) and there is no net flux of energy through the domain boundary, formulation (11) preserves energy, i.e.,
$\partial_{t}\mathcal{E}_{h}=\partial_{t}\mathcal{K}_{h}+\partial_{t}\mathcal{M}_{h}=0.$
4 Fully-discrete weak formulation
If a conservative temporal discrete scheme, for example, the Crank-Nicolson scheme, is applied to the semi-discrete formulation (11) directly, we can obtain a fully-discrete formulation of same conservation and dissipation properties (except that the time derivative of energy is now expressed in a temporally discrete format). The proof is straightforward; it mimics the analyses in Section 3.2. Therefore, the rest of this paper is devoted to a decoupled formulation with computational efficiency in mind.
4.1 Decoupled temporal discretization
Consider a set of time instants,
$\left\{t^{0},t^{\frac{1}{2}},t^{1},t^{1+\frac{1}{2}},t^{2},\cdots\right\},$
where $t^{0}=0$, $\Delta t=t^{k}-t^{k-1}>0$, $k\in\left\{1,2,3,\cdots\right\}$, is a constant time interval, and $t^{k-\frac{1}{2}}=\dfrac{t^{k-1}+t^{k}}{2}$. We use a superscript to denote the evaluation of a variable at a particular time instant. For example, $\bm{u}^{k}_{h}:=\bm{u}_{h}\left(\bm{x},t^{k}\right)$.
A fully discrete, temporally decoupled formulation of (11) (hereinafter referred to as the decoupled formulation) is expressed as follows. Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{\omega}_{h}^{0},\bm{H}^{\frac{1}{2}}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)\times C\left(\Omega\right)$, for time step index $k=1,2,3,\cdots$ successively,
(step 1) seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C_{\widehat{\bm{\omega}}}^{\parallel}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)$, such that, $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)$,
(22a) $\left\langle\dfrac{\bm{u}^{k}_{h}-\bm{u}^{k-1}_{h}}{\Delta t},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\dfrac{\bm{\omega}^{k-1}_{h}+\bm{\omega}^{k}_{h}}{2},\dfrac{\bm{u}^{k-1}_{h}+\bm{u}^{k}_{h}}{2},\bm{v}_{h}\right)$ +Rf1

### v2

@@
 Abstract
-In the framework of a mixed finite element method, a structure-preserving formulation for incompressible MHD equations with general boundary conditions is proposed.
+In the framework of a mixed finite element method, a structure-preserving formulation for incompressible magnetohydrodynamic (MHD) equations with general boundary conditions is proposed.
 A leapfrog-type temporal scheme fully decouples the fluid part from the Maxwell part by means of staggered discrete time sequences and, in doing so, partially linearizes the system.
 Conservation and dissipation properties of the formulation before and after the decoupling are analyzed.
-We demonstrate optimal spatial and second-order temporal error convergence and conservation and dissipation properties of the proposed method using manufactured solutions, and apply it to the benchmark Orszag-Tang and lid-driven cavity test cases.
+We demonstrate optimal spatial and second-order temporal accuracy, as well as conservation and dissipation properties, of the proposed method using manufactured solutions, and apply it to the benchmark Orszag-Tang and lid-driven cavity cases.
 keywords:
-MHD , mixed finite element method , structure-preserving discretization , general boundary conditions
+MHD , structure-preserving discretization , decoupled discretization , mixed finite element method , general boundary conditions
 journal: arXiv
@@
 […]
-[1]organization=School of Mathematics and Computing Science, Guilin University of Electronic Technology, city=Guilin, country=China
+[1]organization=Guilin University of Electronic Technology, School of Mathematics and Computing Science, city=Guilin, country=China
 […]
@@
 The material parameters involved are the fluid density $\rho$, the dynamic viscosity $\tilde{\mu}$, the electric conductivity $\sigma$, and the magnetic permeability $\mu$.
-The incompressible MHD equations, (1), are a coupled system of the incompressible Navier-Stokes equations and the Maxwell’s equations.
-The two evolution equations, (1a) and (1c), stand for the conservation of momentum and the Faraday’s law of induction, respectively.
+The incompressible MHD equations, (1), are a coupled system of the incompressible Navier-Stokes equations and Maxwell’s equations.
+The two evolution equations, (1a) and (1c), stand for the conservation of momentum and Faraday’s law of induction, respectively.
 Ohm’s law and Ampere’s law are (1d) and (1e), respectively.
@@
 The divergence-free constraint on $\bm{u}$, (1b), represents mass conservation of the constant mass density flow.
-It is seen that, as $\bm{B}^{0}:=\bm{B}(\bm{x},0)$ is divergence free, (1c) ensures
+It is seen that, as $\bm{B}^{0}:=\bm{B}(\bm{x},0)$ is divergence-free, (1c) ensures
 $\nabla\cdot\bm{B}=0\quad\mathrm{in}\ \Omega\times(0,T],$
@@
 And, from (1e), we know $\nabla\cdot\bm{j}=0$ which stands for conservation of charge.
-The coupling between the flow field and the electromagnetic field is through the Lorentz force, $\bm{j}\times\bm{B}$ in (1a), and the electromotive force, $\bm{u}\times\bm{B}$ in (1d).
+The coupling between the flow field and the electromagnetic field is achieved through the terms $\bm{j}\times\bm{B}$ in (1a) and $\bm{u}\times\bm{B}$ in (1d).
 Let $L$, $U$, and $B$ denote characteristic quantities of length, velocity, and magnetic flux density, respectively. We scale the space-time domain and variables by
@@
 $\Delta\leftarrow L^{2}\Delta,\quad\partial_{t}\leftarrow\dfrac{L}{U}\partial_{t},\quad\bm{u}\cdot\nabla\leftarrow\dfrac{L}{U}\left(\bm{u}\cdot\nabla\right).$
-Note that we do not introduce new notations for the scaled objects and, after this point, they always refer to the scaled dimensionless ones.
-Now (1) can be written in a dimensionless form, i.e.,
+Note that we do not introduce new notation for the scaled objects, and, after this point, they always refer to the scaled dimensionless ones.
+Now (1) can be written in a dimensionless form,
 (2a) $\partial_{t}\bm{u}+\bm{\omega}\times\bm{u}+\mathrm{R}_{f}^{-1}\nabla\times\bm{\omega}-\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}+\nabla P$ $=\bm{f},$ in $\qquad\text{in }\Omega\times(0,T],$
@@
 This is the case in vector calculus.
-In differential forms, they appear as a pair of differently oriented dual forms connected by the Hodge $\star$ operator, i.e., $\bm{B}=\star\bm{H}$.
+In differential forms, they appear as a pair of differently oriented dual forms connected by the Hodge $\star$ operator, i.e.
+$\bm{B}=\star\bm{H}$.
 The three dimensionless parameters are the fluid Reynolds number $\mathrm{R}_{f}=\dfrac{\rho UL}{\tilde{\mu}}=\dfrac{UL}{\nu}$ (with $\nu=\dfrac{\tilde{\mu}}{\rho}$ being the kinematic viscosity), the magnetic Reynolds number $\mathrm{R}_{m}=\mu\sigma UL$, and the coupling number $\mathsf{c}=\mathrm{A}_{l}^{-2}$ ($\mathrm{A}_{l}=\dfrac{U\sqrt{\rho\mu}}{B}=\dfrac{U}{U_{\mathrm{A}}}$ is the Alfvén number and $U_{\mathrm{A}}=\dfrac{B}{\sqrt{\rho\mu}}$ is the Alfvén speed).
@@
 For example, see [3, (2.2a)].
+One can see that, from (3),
+$\partial_{t}\bm{H}=\partial_{t}\bm{B}=-\nabla\times\bm{E},$
+and due to this, we will have that
+(4) $\partial_{t}\nabla\cdot\bm{H}=-\nabla\cdot\nabla\times\bm{E}=0,$
+which leads to Gauss’s law for magnetism if $\bm{H}^{0}$ is divergence-free.
 To close the problem, (2) is complemented with initial conditions $\bm{u}^{0},\bm{H}^{0}$ and boundary conditions. In this work, we consider the following general boundary condition configuration:
@@
     For the fluid part, there are two pairs of boundary conditions,
-    (4) {$&P=\widehat{P}\quad&&\text{on }\Gamma_{\widehat{P}}\times\left(0,T\right]\\ &\bm{u}\cdot\bm{n}=\widehat{u}\quad&&\text{on }\Gamma_{\widehat{u}}\times\left(0,T\right]\end{aligned}\right.\quad\text{and}\quad\left\{\begin{aligned} &\bm{u}\times\bm{n}=\widehat{\bm{u}}\quad&&\text{on }\Gamma_{\widehat{\bm{u}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{\omega}\times\bm{n}\right)=\widehat{\bm{\omega}}\quad&&\text{on }\Gamma_{\widehat{\bm{\omega}}}\times\left(0,T\right]\end{aligned}\right.,$
-    where $\bm{n}$ is the outward unit normal vector, $\left\{\Gamma_{\widehat{P}},\Gamma_{\widehat{u}}\right\}$ and $\left\{\Gamma_{\widehat{\bm{u}}},\Gamma_{\widehat{\bm{\omega}}}\right\}$ are two partitions of the boundary, such that $\overline{\Gamma}_{\widehat{P}}\cup\overline{\Gamma}_{\widehat{u}}=\partial\Omega$, $\Gamma_{\widehat{P}}\cap\Gamma_{\widehat{u}}=\emptyset$ and $\overline{\Gamma}_{\widehat{\bm{u}}}\cup\overline{\Gamma}_{\widehat{\bm{\omega}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{u}}}\cap\Gamma_{\widehat{\bm{\omega}}}=\emptyset$.
+    (5) {$&P=\widehat{P}\quad&&\text{on }\Gamma_{\widehat{P}}\times\left(0,T\right]\\ &\bm{u}\cdot\bm{n}=\hat{u}\quad&&\text{on }\Gamma_{\hat{u}}\times\left(0,T\right]\end{aligned}\right.\quad\text{and}\quad\left\{\begin{aligned} &\bm{u}\times\bm{n}=\widehat{\bm{u}}\quad&&\text{on }\Gamma_{\widehat{\bm{u}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{\omega}\times\bm{n}\right)=\widehat{\bm{\omega}}\quad&&\text{on }\Gamma_{\widehat{\bm{\omega}}}\times\left(0,T\right]\end{aligned}\right.,$
+    where $\bm{n}$ is the outward unit normal vector, $\left\{\Gamma_{\widehat{P}},\Gamma_{\hat{u}}\right\}$ and $\left\{\Gamma_{\widehat{\bm{u}}},\Gamma_{\widehat{\bm{\omega}}}\right\}$ are two partitions of the boundary, such that $\overline{\Gamma}_{\widehat{P}}\cup\overline{\Gamma}_{\hat{u}}=\partial\Omega$, $\Gamma_{\widehat{P}}\cap\Gamma_{\hat{u}}=\emptyset$ and $\overline{\Gamma}_{\widehat{\bm{u}}}\cup\overline{\Gamma}_{\widehat{\bm{\omega}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{u}}}\cap\Gamma_{\widehat{\bm{\omega}}}=\emptyset$.
   • 2.
@@
 […]
-    on $\partial\Omega\times\left(0,T\right]$. Also see [4, Section 2.3].
+    on $\partial\Omega\times\left(0,T\right]$. Also, see [4, Section 2.3].
 For example, the boundary conditions considered in the work of Hu et al., see (7) of [5], are a particular case of this general configuration, i.e.
-the case when $\Gamma_{\widehat{u}}=\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset$, $\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$.
-It is generally acknowledged that structure-preservation, especially the preservation (either strongly of weakly) of $\nabla\cdot\bm{B}=0$, is an essential property of numerical methods for MHD [6].
+the case when $\Gamma_{\hat{u}}=\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset$, $\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$.
+It is generally acknowledged that structure-preservation, especially the preservation (either strongly or weakly) of $\nabla\cdot\bm{B}=0$, is an essential property of numerical methods for MHD [6].
 Many of these structure-preserving methods can be classified into, for example, potential-based methods [7], divergence-cleaning methods [8, 9], constrained transport methods [10, 11] and Lagrange multiplier (or augmented) methods [12, 13, 14].
-In the current paper neither of those techniques will be applied, but it will be shown that the absence of magnetic monopoles is weakly enforced through the decoupled variational formulation.
-Recently, methods that aim at preserving multiple structures in the mixed finite element setting become increasingly popular.
+In the current paper, none of those techniques will be applied, but it will be shown that the absence of magnetic monopoles is weakly enforced through the decoupled variational formulation.
+Recently, methods that aim at preserving multiple structures in the mixed finite element setting have become increasingly popular.
 See, for example, the work of Hu et al.
-for incompressible MHD that preserves cross- and magnetic-helicity, energy and Gauss law of magnetism [5], the work of Gawlik and Gay-Balmaz for incompressible MHD of a variable fluid density that preserves energy, cross-helicity (when the fluid density is constant) and magnetic-helicity, mass, total squared density, pointwise incompressibility, and Gauss’s law for magnetism [15], the work of Laakmann et al.
+for incompressible MHD that preserves cross- and magnetic-helicity, energy, and Gauss law of magnetism [5], the work of Gawlik and Gay-Balmaz for incompressible MHD of a variable fluid density that preserves energy, cross-helicity (when the fluid density is constant) and magnetic-helicity, mass, total squared density, pointwise incompressibility, and Gauss’s law for magnetism [15], the work of Laakmann et al.
 [16] and references therein.
-These methods preserve several physical quantities of interest, but usually are less computationally efficient because of their large (nonlinear) discrete systems to be solved.
-A compromise is to use a decoupled scheme which breaks the discrete system into several smaller ones, for example see [3, 12, 13, 17, 18].
-These methods in the mixed finite element setting commonly deal with a particular configuration of boundary conditions.
-In this paper, we introduce a structure-preserving mixed finite element discretization for the incompressible MHD under a general boundary condition configuration.
-The method is decoupled in the sense that the two evolution equations are temporally discretized at different time sequences such that they can be solved individually.
+These methods preserve several physical quantities of interest, but are usually less computationally efficient because of their large (nonlinear) discrete systems to be solved.
+A compromise is to use a decoupled scheme that breaks the discrete system into several smaller ones.
+For example, Zhang and Su [3] propose a three-step decoupled scheme by using a first-order semi-implicit Euler scheme with a first-order stabilizer and some carefully designed implicit-explicit treatments for the coupling terms.
+In [13], a four-step decoupled scheme of second-order temporal accuracy is studied.
+It employs the mixed finite element method under a “zero-energy-contribution” idea and a projection method.
+Ma and Huang introduced a vector penalty-projection method, a five-step decoupled scheme that has a mixed finite element spatial discretization and a first-order backward Euler temporal discretization in [17].
+The work [18] reports another five-step decoupled scheme which combines the idea of an auxiliary variable, pressure correction projection method, and some other implicit-explicit techniques for the thermally coupled MHD equations with variable density.
+For a more extended literature study on decoupled schemes in a mixed finite element setting for MHD equations, we refer to [17].
+These methods commonly deal with a particular configuration of boundary conditions.
+In this paper, we introduce a mixed finite element structure-preserving decoupled discretization of a second-order temporal accuracy for the incompressible MHD equations with general boundary conditions.
+This method decouples each full iteration into two steps by means of an implicit mid-point scheme that is applied to the two evolution equations at staggered temporal sequences such that they can be solved individually in a leapfrog approach.
 The outline of the rest of the paper is as follows.
@@
 Numerical tests are presented in Section 5.
-Finally, in Section 6 conclusions are drawn.
+Finally, in Section 6, conclusions are drawn.
 2 Relevant Hilbert spaces and the de Rham complex
@@
 For the sake of completeness, we give a brief introduction to the three-dimensional spaces as a complement to [19, Section 2.1].
-We first introduce the relevant spaces and the de Rham complex in the infinite dimensional setting, and then introduce their finite dimensional counterparts.
-For an extensive introduction of function spaces, see [20] or any other textbook on functional analysis.
-2.1 The infinite dimensional setting
+We first introduce the relevant spaces and the de Rham complex in the infinite-dimensional setting and then introduce their finite-dimensional counterparts.
+For an extensive introduction to function spaces, see [20] or any other textbook on functional analysis.
+2.1 The infinite-dimensional setting
 Let $L^{2}\left(\Omega\right)$ denote the space of square-integrable functions on $\Omega$,
@@
 $H\left(\mathrm{div};\Omega\right)$ :=$:=\left\{\bm{\phi}\left|\bm{\phi}\in\left[L^{2}\left(\Omega\right)\right]^{3},\nabla\cdot\bm{\phi}\in L^{2}\left(\Omega\right)\right.\right\}.$
-They form the well known de Rham Hilbert complex [21, 22],
-(7) $\mathbb{R}\hookrightarrow H^{1}\left(\Omega\right)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}H(\mathrm{curl};\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}H(\mathrm{div};\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}L^{2}(\Omega)\to 0.$
+They form the well-known de Rham Hilbert complex [21, 22],
+$\mathbb{R}\hookrightarrow H^{1}\left(\Omega\right)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}H(\mathrm{curl};\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}H(\mathrm{div};\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}L^{2}(\Omega)\to 0.$
 […]
@@
 $\mathcal{T}H\left(\mathrm{div};\Omega,\Gamma\right):=\left\{\mathcal{T}\bm{\phi}\left|\bm{\phi}\in H(\mathrm{div};\Omega)\right.\right\}.$
-Note that, at the continuous level, although $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ are different, see (8), trace spaces $\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right)$ and $\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right)$ refer to the same tangential space.
-2.2 The finite dimensional setting
-Assume that we are given finite dimensional spaces $G(\Omega)\subset H^{1}(\Omega)$, $C(\Omega)\subset H(\mathrm{curl};\Omega)$, $D(\Omega)\subset H(\mathrm{div};\Omega)$, $S(\Omega)\subset L^{2}(\Omega)$, which form a discretization of the continuous de Rham Hilbert complex,
-(9) $\mathbb{R}\hookrightarrow G(\Omega)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}C(\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}D(\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}S(\Omega)\to 0.$
-In particular, this means that if the complex in (7) is exact, then so is the one in (9). We will also consider the following subspaces with boundary conditions, where $\Gamma\subset\partial\Omega$,
+Note that, at the continuous level, $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ are different, see (8), and trace spaces $\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right)$ and $\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right)$ are in a duality, see [23].
+2.2 The finite-dimensional setting
+Assume that we are given finite-dimensional spaces $G(\Omega)\subset H^{1}(\Omega)$, $C(\Omega)\subset H(\mathrm{curl};\Omega)$, $D(\Omega)\subset H(\mathrm{div};\Omega)$, $S(\Omega)\subset L^{2}(\Omega)$ which form a discretization of the continuous de Rham Hilbert complex,
+(9) $\mathbb{R}\hookrightarrow G(\Omega)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}C(\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}D(\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}S(\Omega)\to 0,$
+which is also exact in the contractible domain $Omega$. We will also consider the following subspaces with boundary conditions, where $\Gamma\subset\partial\Omega$,
 $G_{\widehat{\psi}}(\Omega,\Gamma):=\left\{\psi_{h}\left|\psi_{h}\in G(\Omega),\ \mathcal{T}\psi_{h}=\widehat{\psi}\ \text{on}\ \Gamma\right.\right\},$
@@
 $D_{0}(\Omega,\Gamma):=\left\{\bm{\phi}_{h}\left|\bm{\phi}_{h}\in D(\Omega),\ \mathcal{T}\bm{\phi}_{h}=0\ \text{on}\ \Gamma\right.\right\}.$
-Moreover, the finite dimensional trace spaces will be denoted by $\mathcal{T}G(\Omega,\Gamma)$, $\mathcal{T}C(\Omega,\Gamma)$, $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$, $\mathcal{T}D(\Omega,\Gamma)$.
-Unlike their continuous counterparts, finite dimensional trace spaces $\mathcal{T}C(\Omega,\Gamma)$ and $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$ usually are different.
+Moreover, the finite-dimensional trace spaces will be denoted by $\mathcal{T}G(\Omega,\Gamma)$, $\mathcal{T}C(\Omega,\Gamma)$, $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$, $\mathcal{T}D(\Omega,\Gamma)$.
+Unlike their continuous counterparts, finite-dimensional trace spaces $\mathcal{T}C(\Omega,\Gamma)$ and $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$ usually are different.
 We define a corresponding trilinear form,
@@
 which is skew-symmetric with respect to any two of the three entries.
-Note that, in the infinite dimensional setting, the $L^{2}$-boundedness of $\left\langle{\bm{\alpha}\times\bm{\beta}},{\bm{\gamma}}\right\rangle_{\Omega}$ is not implied even if $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are $L^{2}$ vector fields.
-However, in the finite dimensional setting where the spaces $C(\Omega)$ and $D(\Omega)$ consist of piecewise-polynomial finite element vector fields, it is in fact $L^{2}$-bounded.
+Note that, in the infinite-dimensional setting, the $L^{2}$-boundedness of $\left\langle{\bm{\alpha}\times\bm{\beta}},{\bm{\gamma}}\right\rangle_{\Omega}$ is not implied even if $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are $L^{2}$ vector fields.
+However, in the finite-dimensional setting where the spaces $C(\Omega)$ and $D(\Omega)$ consist of piecewise-polynomial finite element vector fields, it is in fact $L^{2}$-bounded.
 3 Spatially-discrete formulation and its conservation properties
-In this section, we present a spatially-discrete weak formulation of (2) and analyze its conservation properties.
-As a remark, for neatness, we only discuss the three-dimensional formulation and its temporal discretization in this paper.
+In this section, we present a spatially discrete weak formulation of (2) and analyze its conservation properties.
+As a remark, for the sake of neatness, we discuss only the three-dimensional formulation and its temporal discretization in this paper.
 Obtaining their two-dimensional versions is straightforward.
@@
 To increase the notational clarity, we omit the part for the temporal domain.
-The proposed semi-discrete weak formulation of (2) in $\mathbb{R}^{3}$ is written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, seek $\left(\bm{u}_{h},\bm{\omega}_{h},P_{h},\bm{H}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}_{\perp}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
+The proposed semi-discrete weak formulation of (2) in $\mathbb{R}^{3}$ is written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, seek $\left(\bm{u}_{h},\bm{\omega}_{h},P_{h},\bm{H}_{h}\right)\in D_{\hat{u}}(\Omega,\Gamma_{\hat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\hat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
 (11a) $\left\langle\partial_{t}\bm{u}_{h},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\bm{\omega}_{h},\bm{u}_{h},\bm{v}_{h}\right)+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\bm{\omega}_{h}},{\bm{v}_{h}}\right\rangle_{\Omega}\qquad$
@@
 (11d) $\left\langle{\partial_{t}\bm{H}_{h}},{\bm{b}_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\bm{b}_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\bm{b}_{h}\right)$ $=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{b}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}.$
-Note that this assumes that the supplied boundary conditions are exactly representable as members of the chosen finite dimensional spaces; more generally, one can always project the boundary conditions into the appropriate finite dimensional trace spaces.
+Note that this assumes that the supplied boundary conditions are exactly representable as members of the chosen finite-dimensional spaces; more generally, one can always project the boundary conditions into the appropriate finite-dimensional trace spaces.
 See (5) and (6) for the general boundary setting.
@@
 3.2.3 Weak conservation of Gauss’s law for magnetism
+One of the most common forms of the MHD equations is the conservative form (or divergence form); see, for example, [24].
+The discretization of the conservative form poses several challenges.
+One of these challenges is the involution constraint $\nabla\cdot\bm{H}=0$ (or $\nabla\cdot\bm{B}=0$), which has been widely discussed in the literature, e.g.
+[8, 10, 25, 26, 27, 28].
+The main root of this challenge is that this formulation is only valid if the involution constraint is satisfied [24, 26].
+Moreover, in conservation form, the involution constraint is not an implication of these equations and, therefore, must be enforced.
+On the other hand, the formulation we employ in this work is self-consistent with regard to the involution constraint.
+By self-consistent, we mean the following: The involution constraint must only be imposed on the initial condition.
+Once this condition is satisfied by the initial condition, the evolution equation enforces it at all time instants.
+As long as numerical discretization preserves a discrete de Rham complex, this equation will automatically be satisfied and the involution constraint will be enforced if the initial condition is divergence-free; see (4).
+In a different context, it has been reported that the formulation leads to better results than the conservation form [26], when not enforcing the involution constraint.
+More recently, [24], a similar weak formulation has been proposed.
+In that work, as in this, the weak divergence-free constraint is preserved throughout the whole evolution.
+As seen in the next section (and in the numerical results), this does not lead to a loss of energy conservation, in contrast to approaches for the conservative form that do not strongly satisfy the divergence-free constraint.
 Since (11d) is valid for all $\bm{b}_{h}\in C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$, let us choose $\bm{b}_{h}=\nabla q_{h}$ for an arbitrary $q_{h}\in G_{0}(\Omega,\Gamma_{\widehat{\bm{H}}})$.
@@
 (12) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\nabla q_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\nabla q_{h}\right)=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\nabla q_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\;.$
-The second and third terms on the left hand side are zero because we strongly satisfy ×$\nabla\times\nabla q_{h}=0$ since our discrete function spaces constitute a finite dimensional de Rham complex.
+The second and third terms on the left-hand side are zero because we strongly satisfy ×$\nabla\times\nabla q_{h}=0$ since our discrete function spaces constitute a finite-dimensional de Rham complex.
 We are left with the following equation,
@@
 Suppose
-(17) $\widehat{u}=0\;,\quad\Gamma_{\widehat{\bm{u}}}=\Gamma_{\widehat{\bm{E}}}=\partial\Omega\;,\quad\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset\;.$
+(17) $\hat{u}=0\;,\quad\Gamma_{\widehat{\bm{u}}}=\Gamma_{\widehat{\bm{E}}}=\partial\Omega\;,\quad\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset\;.$
 By replacing $\bm{v}_{h}$ in (11a) with $\bm{u}_{h}$, we can obtain
@@
 The two nonlinear terms have canceled each other due to the skew-symmetry of the trilinear form (10); the exchange of energy between $\mathcal{K}_{h}$ and $\mathcal{M}_{h}$ is exactly captured.
-In (21), (i) represents amount of mechanical work done on the system per unit time, (ii) and (iii) are dissipation rates due to viscosity and electric resistance, respectively, and terms (iv) - (vi) are the net flux of energy through the domain boundary.
+In (21), (i) represents the amount of mechanical work done on the system per unit time, (ii) and (iii) are dissipation rates due to viscosity and electric resistance, respectively, and terms (iv) - (vi) are the net flux of energy through the domain boundary.
 Clearly, if $\bm{f}=\bm{0}$, when the MHD flow is ideal (i.e.
-$\mathrm{R}_{f}=\mathrm{R}_{m}=\infty$) and there is no net flux of energy through the domain boundary, formulation (11) preserves energy, i.e.,
+$\mathrm{R}_{f}=\mathrm{R}_{m}=\infty$) and there is no net flux of energy through the domain boundary, formulation (11) preserves energy, i.e.
 $\partial_{t}\mathcal{E}_{h}=\partial_{t}\mathcal{K}_{h}+\partial_{t}\mathcal{M}_{h}=0.$
-4 Fully-discrete weak formulation
-If a conservative temporal discrete scheme, for example, the Crank-Nicolson scheme, is applied to the semi-discrete formulation (11) directly, we can obtain a fully-discrete formulation of same conservation and dissipation properties (except that the time derivative of energy is now expressed in a temporally discrete format).
+4 Fully discrete weak formulation
+If a conservative temporal discrete scheme, for example, the Crank-Nicolson scheme, is applied to the semi-discrete formulation (11) directly, we can obtain a fully discrete formulation of the same conservation and dissipation properties (except that the time derivative of energy is now expressed in a temporally discrete format).
+See B.
 The proof is straightforward; it mimics the analyses in Section 3.2.
@@
 4.1 Decoupled temporal discretization
-Consider a set of time instants,
-$\left\{t^{0},t^{\frac{1}{2}},t^{1},t^{1+\frac{1}{2}},t^{2},\cdots\right\},$
+Consider a set of time instants
+(22) $\left\{t^{0},t^{\frac{1}{2}},t^{1},t^{1+\frac{1}{2}},t^{2},\cdots\right\},$
 where $t^{0}=0$, $\Delta t=t^{k}-t^{k-1}>0$, $k\in\left\{1,2,3,\cdots\right\}$, is a constant time interval, and $t^{k-\frac{1}{2}}=\dfrac{t^{k-1}+t^{k}}{2}$.
@@
 A fully discrete, temporally decoupled formulation of (11) (hereinafter referred to as the decoupled formulation) is expressed as follows.
-Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{\omega}_{h}^{0},\bm{H}^{\frac{1}{2}}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)\times C\left(\Omega\right)$, for time step index $k=1,2,3,\cdots$ successively,
-(step 1) seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C_{\widehat{\bm{\omega}}}^{\parallel}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)$, such that, $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)$,
-(22a) $\left\langle\dfrac{\bm{u}^{k}_{h}-\bm{u}^{k-1}_{h}}{\Delta t},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\dfrac{\bm{\omega}^{k-1}_{h}+\bm{\omega}^{k}_{h}}{2},\dfrac{\bm{u}^{k-1}_{h}+\bm{u}^{k}_{h}}{2},\bm{v}_{h}\right)$ +Rf1
+Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{\omega}_{h}^{0},\bm{H}^{\frac{1}{2}}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)\times C\left(\Omega\right)$, for time-step index $k=1,2,3,\cdots$ successively,
+(step 1) seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h}\right)\in D_{\hat{u}}(\Omega,\Gamma_{\hat{u}})\times C_{\widehat{\bm{\omega}}}^{\parallel}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)$, such that, $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\hat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)$,
+(23a) $\left\langle\dfrac{\bm{u}^{k}_{h}-\bm{u}^{k-1}_{h}}{\Delta t},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\dfrac{\bm{\omega}^{k-1}_{h}+\bm{\omega}^{k}_{h}}{2},\dfrac{\bm{u}^{k-1}_{h}+\bm{u}^{k}_{h}}{2},\bm{v}_{h}\right)$ $+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\dfrac{\bm{\omega}^{k-1}_{h}+\bm{\omega}^{k}_{h}}{2}},{\bm{v}_{h}}\right\rangle_{\Omega}$
+$-\mathsf{c}\ a\left(\nabla\times\bm{H}^{k-\frac{1}{2}}_{h},\bm{H}^{k-\frac{1}{2}}_{h},\bm{v}_{h}\right)-\left\langle{P^{k-\frac{1}{2}}_{h}},{\nabla\cdot\bm{v}_{h}}\right\rangle_{\Omega}$ $=\left\langle{\bm{f}^{k-\frac{1}{2}}},{\bm{v}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}^{k-\frac{1}{2}}},{\mathcal{T}\bm{v}_{h}}\right\rangle_{\Gamma_{\widehat{P}}},$
+(23b) $-\left\langle{\bm{u}^{k}_{h}},{\nabla\times\bm{w}_{h}}\right\rangle_{\Omega}+\left\langle{\bm{\omega}^{k}_{h}},{\bm{w}_{h}}\right\rangle_{\Omega}$ $=-\left\langle{\widehat{\bm{u}}^{k}},{\mathcal{T}_{\parallel}\bm{w}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{u}}}},$
+(23c) $\left\langle{\nabla\cdot\bm{u}^{k}_{h}},{q_{h}}\right\rangle_{\Omega}$ $=0,$
+(step 2) seek $\bm{H}^{k+\frac{1}{2}}_{h}\in C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that, $\forall\bm{b}_{h}\in C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
+(24) $\left\langle{\dfrac{\bm{H}^{k+\frac{1}{2}}_{h}-\bm{H}^{k-\frac{1}{2}}_{h}}{\Delta t}},{\bm{b}_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\dfrac{\bm{H}^{k-\frac{1}{2}}_{h}+\bm{H}^{k+\frac{1}{2}}_{h}}{2}},{\nabla\times\bm{b}_{h}}\right\rangle_{\Omega}$
+$-\ a\left(\bm{u}^{k}_{h},\dfrac{\bm{H}^{k-\frac{1}{2}}_{h}+\bm{H}^{k+\frac{1}{2}}_{h}}{2},\nabla\times\bm{b}_{h}\right)$ $=\left\langle{\widehat{\bm{E}}^{k}},{\mathcal{T}_{\parallel}\bm{b}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}.$
+Note that the decoupled formulation takes $\bm{H}_{h}^{\frac{1}{2}}$, instead of $\bm{H}_{h}^{0}$, as an initial condition.
+Thus, to initiate the iterations, we need to compute $\bm{H}_{h}^{\frac{1}{2}}$ by applying, for example, an explicit Euler scheme or the Crank-Nicolson scheme as discussed in Section 4 to (11) at the half-time-step from $t^{0}$ to $t^{\frac{1}{2}}$.
+An illustration of the overall temporal scheme for the decoupled formulation is shown in Fig.
+1.
+Figure 1: An illustration of the leapfrog temporal scheme for the decoupled formulation.
+The time-step sequence is $\hat{s}^{0}\to S^{1}\to\hat{S}^{1}\to S^{2}\to\hat{S}^{2}\to\cdots$.
+Steps $S^{1},S^{2},S^{3},\cdots$ refer to the step 1, see (23), of the decoupled formulation, and steps $\hat{S}^{1},\hat{S}^{2},\hat{S}^{3},\cdots$ refer to the step 2, see (24), of the decoupled formulation.
+The pre-step $\hat{s}^{0}$ computes $\bm{H}_{h}^{\frac{1}{2}}$.
+A similar scheme has been used in a dual-field discretization of incompressible Navier-Stokes equations, see [29, Fig.
+1].
+This leapfrog-type temporal scheme is analogous to the symplectic Störmer–Verlet leapfrog integrator [30] in Hamiltonian mechanics.
+Seeing that the decoupled formulation has linearized the Maxwell part, see (24), it does not fully linearize the Navier-Stokes part.
+More specifically, the second term of (23a), namely, the fluid convection term, is yet to be left non-linear.
+Thus, a nonlinear system solver, for example, the Picard method or the Newton-Raphson method, still needs to be employed.
+4.2 Conservation properties
+Clearly, the formulation (23) strongly preserves conservation of mass, i.e., $\nabla\cdot\bm{u}_{h}^{k}=0$, and conservation of charge, i.e., $\nabla\cdot\bm{j}_{h}^{k+\frac{1}{2}}=0$, pointwise, and weakly preserves Gauss’s law for magnetism, for same reasons as in Section 3.2.1, Section 3.2.2 and Section 3.2.3, respectively; the temporal discretization scheme maintains these conservation properties at the fully discrete level.
+For energy conservation, it has been shown that, for the semi-discrete formulation, a key to deriving it is the canceling of the two nonlinear terms, see (19) - (21).
+We now study whether a similar canceling happens for the decoupled formulation.
+And, to simplify the notations, we use a periodic domain, i.e., whose $\partial\Omega=\emptyset$ or vanishing boundary conditions, i.e.
+$\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$ on $\partial\Omega\times(0,T]$.
+Throughout the paper, we use the midpoint rule, i.e., for example,
+$\bm{u}_{h}^{k-\frac{1}{2}}:=\dfrac{\bm{u}_{h}^{k-1}+\bm{u}_{h}^{k}}{2}.$
+If we replace $\bm{v}_{h}$ in (23a) with $\bm{u}_{h}^{k-\frac{1}{2}}$, by performing similar analyses as those in Section 3.2.4, we can find that
+$\left\langle\frac{\bm{u}^{k}_{h}-\bm{u}^{k-1}_{h}}{\Delta t},\frac{\bm{u}^{k-1}_{h}+\bm{u}^{k}_{h}}{2}\right\rangle_{\Omega}+\mathrm{R}_{f}^{-1}\left\langle{\bm{\omega}_{h}^{k-\frac{1}{2}}},{\bm{\omega}_{h}^{k-\frac{1}{2}}}\right\rangle_{\Omega}-\mathsf{c}\ a\left(\bm{j}^{k-\frac{1}{2}}_{h},\bm{H}^{k-\frac{1}{2}}_{h},\bm{u}^{k-\frac{1}{2}}_{h}\right)=\left\langle{\bm{f}^{k-\frac{1}{2}}},{\bm{u}^{k-\frac{1}{2}}_{h}}\right\rangle_{\Omega},$
+where $\bm{j}_{h}^{k-\frac{1}{2}}=\nabla\times\bm{H}_{h}^{k-\frac{1}{2}}$.
+This leads to, if $\mathcal{K}^{k}:=\dfrac{1}{2}\left\langle{\bm{u}^{k}},{\bm{u}^{k}_{h}}\right\rangle_{\Omega}$ denotes the discrete kinetic energy,
+(25) $\frac{\mathcal{K}_{h}^{k}-\mathcal{K}_{h}^{k-1}}{\Delta t}=\mathcal{F}^{k-\frac{1}{2}}-\mathrm{R}_{f}^{-1}\mathcal{S}_{h}^{k-\frac{1}{2}}+\mathsf{c}\mathcal{A}_{h}^{k-\frac{1}{2}},$
+where $\mathcal{F}^{k-\frac{1}{2}}:=\left\langle{\bm{f}^{k-\frac{1}{2}}},{\bm{u}^{k-\frac{1}{2}}_{h}}\right\rangle_{\Omega}$, $\mathcal{S}_{h}^{k-\frac{1}{2}}:=\left\langle{\bm{\omega}_{h}^{k-\frac{1}{2}}},{\bm{\omega}_{h}^{k-\frac{1}{2}}}\right\rangle_{\Omega}$ and $\mathcal{A}_{h}^{k-\frac{1}{2}}:=a\left(\bm{j}^{k-\frac{1}{2}}_{h},\bm{H}^{k-\frac{1}{2}}_{h},\bm{u}^{k-\frac{1}{2}}_{h}\right)$.
+Similarly, by replacing $\bm{b}_{h}$ in (24) with $\bm{H}^{k}_{h}$, we can obtain
+(26) $\frac{\mathcal{M}_{h}^{k+\frac{1}{2}}-\mathcal{M}_{h}^{k-\frac{1}{2}}}{\Delta t}=-\mathsf{c}\mathrm{R}_{m}^{-1}\mathcal{J}_{h}^{k}-\mathsf{c}\mathcal{A}^{k}_{h},$
+where $\mathcal{M}_{h}^{k+\frac{1}{2}}=\dfrac{\mathsf{c}}{2}\left\langle{\bm{H}^{k+\frac{1}{2}}},{\bm{H}^{k+\frac{1}{2}}}\right\rangle_{\Omega}$ denotes the discrete magnetic energy, and $\mathcal{J}_{h}^{k}:=\left\langle{\bm{j}^{k}_{h}},{\bm{j}^{k}_{h}}\right\rangle_{\Omega}$.
+Repeating this analysis for the time-step indexed $(k-1)$ gives
+(27) $\frac{\mathcal{M}_{h}^{k-\frac{1}{2}}-\mathcal{M}_{h}^{k-\frac{3}{2}}}{\Delta t}=-\mathsf{c}\mathrm{R}_{m}^{-1}\mathcal{J}_{h}^{k-1}-\mathsf{c}\mathcal{A}^{k-1}_{h}.$
+(26) and (27) together shows
+(28) ~$\frac{\tilde{\mathcal{M}}_{h}^{k}-\tilde{\mathcal{M}}_{h}^{k-1}}{\Delta t}=-\mathsf{c}\mathrm{R}_{m}^{-1}\tilde{\mathcal{J}}_{h}^{k-\frac{1}{2}}-\mathsf{c}\tilde{\mathcal{A}}^{k-\frac{1}{2}}_{h},$
+where we have used the tilde sign to denote that it is an average of two integral quantities, i.e., for example,
+(29) ~$\tilde{\mathcal{M}}_{h}^{k}:=\dfrac{\mathcal{M}_{h}^{k-\frac{1}{2}}+\mathcal{M}_{h}^{k+\frac{1}{2}}}{2}$
+which differs from ${\mathcal{M}}_{h}^{k}=\dfrac{\mathsf{c}}{2}\left\langle{\bm{H}^{k}},{\bm{H}^{k}}\right\rangle_{\Omega}.$ With (25) and (28), we can conclude that the decoupled formulation dissipates a discrete energy defined as
+(30) ~$\tilde{\mathcal{E}}_{h}^{k}:={\mathcal{K}}_{h}^{k}+\tilde{\mathcal{M}}_{h}^{k}$
+with a discrete rate
+(31) ~$\frac{\tilde{\mathcal{E}}_{h}^{k}-\tilde{\mathcal{E}}_{h}^{k-1}}{\Delta t}=\mathcal{F}^{k-\frac{1}{2}}-\mathrm{R}_{f}^{-1}\mathcal{S}_{h}^{k-\frac{1}{2}}-\mathsf{c}\mathrm{R}_{m}^{-1}\tilde{\mathcal{J}}_{h}^{k-\frac{1}{2}}+\mathsf{c}\left(\mathcal{A}_{h}^{k-\frac{1}{2}}-\tilde{\mathcal{A}}^{k-\frac{1}{2}}_{h}\right).$
+It is not guarantee that the formulation will dissipate as the last term is undefined; when $\bm{f}=\bm{0}$ and the flow is ideal, we have
+~$\frac{\tilde{\mathcal{E}}_{h}^{k}-\tilde{\mathcal{E}}_{h}^{k-1}}{\Delta t}=\mathsf{c}\left(\mathcal{A}_{h}^{k-\frac{1}{2}}-\tilde{\mathcal{A}}^{k-\frac{1}{2}}_{h}\right).$
+As generally $\mathcal{A}_{h}^{k-\frac{1}{2}}-\tilde{\mathcal{A}}^{k-\frac{1}{2}}\neq 0$, the decoupled formulation does not conserve the discrete energy (30).
+Exact preservation of the discrete energy would require a fully implicit scheme coupling the fluid and Maxwell parts.
+This can be achieved, for instance, using the midpoint rule as the Hamiltonian is quadratic in the velocity and magnetic field and the midpoint rule preserves quadratic invariants [31].
+Many integration strategies have been proposed in recent years to ensure energy preservation without the need to solve nonlinear systems.
+These schemes rely on the idea of energy quadratization via the introduction of a scalar auxiliary variable and lead to linearly implicit schemes [32].
+The scalar auxiliary variable approach combines the implicit midpoint scheme with a Störmer-Verlet integrator to achieve second-order accuracy in time and energy conservation.
+For the MHD equations, the nonlinear coupling between fluid and electromagnetic systems makes it difficult to achieve second-order accuracy and energy conservation without the need for a fully implicit nonlinear solver.
+In short, the decoupled formulation breaks the nonlinear MHD formulation into two evolution equations representing the Navier-Stokes part and the Maxwell part, respectively, partially linearizes the former and fully linearizes the latter.
+It keeps strong conservation of mass, strong conservation of charge, and weak conservation of Gauss’s law for magnetism while sacrificing conservation of energy.
+More importantly, it allows to keep second-order accuracy in time while confining the nonlinear iterations to the fluid system.
+5 Numerical tests
+In this section, we present some results of numerical tests using the proposed method.
+The test cases are manufactured solution tests, the Orszag-Tang vortex flow, and magnetic lid-driven cavity flow.
+We use the mimetic spectral elements [33, 34] for all tests.
+The degree of the basis functions is denoted by $N$.
+Any other set of finite elements that also satisfies the discrete de Rham Hilbert complex (9) and the $L^{2}$-regularity for the triple products, see Section 2.2 works for the proposed method.
+A combination of Lagrange (continuous Galerkin) elements of degree $N$, the first kind Nédélec $H(\mathrm{curl})$-conforming elements of degree $N$ [35], Raviart–Thomas elements of degree $N$ [36], and discontinuous Galerkin elements of degree $(N-1)$ [37] is a classic candidate [22, 38].
+5.1 Manufactured solution tests
+Two sets of manufactured solutions are employed for convergence tests and conservation and dissipation tests, respectively.
+5.1.1 Convergence tests
+Suppose three-dimensional manufactured solutions
+$\bm{u}=\begin{bmatrix}\cos(x)\sin(y)\sin(z)e^{t}&\sin(x)\cos(y)\sin(z)e^{t}&-2\sin(x)\sin(y)\cos(z)e^{t}\end{bmatrix}^{\mathsf{T}},$
+$\bm{E}=\begin{bmatrix}\cos(x)\sin(y)\sin(z/2)e^{t}&\sin(x/2)\cos(y)\sin(z)e^{t}&-\sin(x)\sin(y)\cos(z)e^{t}\end{bmatrix}^{\mathsf{T}},$
+$P=\cos(x)\cos(y)\cos(z)e^{-t}\quad\text{and}\quad\bm{H}^{0}=\bm{B}^{0}=\bm{0},$
+solve (2).
+Manufactured solutions for the remaining variables can be obtained from $\bm{\omega}=\nabla\times\bm{u}$, $\bm{H}=\bm{B}=\int_{0}^{t}\left(-\nabla\times\bm{E}\right)\mathrm{d}t$, $\bm{j}=\nabla\times\bm{H}$.
+Given $\mathrm{R}_{f}$ and $\mathsf{c}$, we can get the analytical expression of $\bm{f}$ through (2a).
+And if $\mathrm{R}_{m}$ is known, we can find an extra source, denoted by $\bm{e}$, needed to balance the relation
+$\bm{e}=\mathrm{R}_{m}^{-1}\bm{j}-\left(\bm{E}+\bm{u}\times\bm{B}\right).$
+This will lead to an extra term, $\left\langle\bm{e},\nabla\times\bm{b}_{h}\right\rangle_{\Omega}$, in the right hand sides of, for example, (11d) and (24).
+For the convergence tests, the parameters are set to $\mathrm{R}_{f}=1$, $\mathrm{R}_{m}=1$, $\mathsf{c}=1$.
+And the domain is selected to be $\Omega=\left(x,y,z\right)\in\left[0,2\pi\right]^{3}$ whose boundary is partitioned, see (5) and (6), such that
+$\Gamma_{\widehat{P}}=\Gamma_{x}^{-}\cup\Gamma_{y}^{+}\cup\Gamma_{z}^{+},$
+$\Gamma_{\widehat{\bm{u}}}=\Gamma_{x}^{+}\cup\Gamma_{y}^{-}\cup\Gamma_{z}^{+},$
+$\Gamma_{\widehat{\bm{E}}}=\Gamma_{x}^{+}\cup\Gamma_{y}^{+}\cup\Gamma_{z}^{-},$
+where, for example, $\Gamma_{x}^{-}$ represents the face $\left(x,y,z\right)\in 0\times(0,2\pi)\times(0,2\pi)$ and $\Gamma_{x}^{+}$ represents the face $\left(x,y,z\right)\in 2\pi\times(0,2\pi)\times(0,2\pi)$.
+Let $K$ and $\kappa$ be two positive integers.
+We generate a mesh of $K^{3}$ uniform cubic elements in $\Omega$, and $h=\frac{2\pi}{K}$ is the size, i.e.
+the edge length, of elements.
+On this mesh, given initial conditions, boundary conditions, $\bm{f}$ and $\bm{e}$ according to the manufactured solutions, the decoupled formulation, (23) and (24), is solved with a constant time-step interval $\Delta t=\frac{1}{\kappa}$.
+Errors between simulation and manufactured solutions are measured at $t=1$, i.e.
+when time-step index $k=\kappa$.
+Results in Fig.
+2 show that the decoupled temporal discretization has a second-order accuracy for both evolution equations.
+In Fig.
+3, optimal spatial convergence rates are observed for all variables.
+Results supporting that the formulation weakly preserves Gauss’ law for magnetism are presented in Fig.
+4.
+Figure 2: Results of $\left\|\bm{u}^{k}_{h}\right\|_{H(\mathrm{div})\text{-error}}$ and $\left\|\bm{H}^{k+\frac{1}{2}}_{h}\right\|_{H(\mathrm{curl})\text{-error}}$ for the temporal convergence tests at $N=3$, $K=16$, $\Delta t\in\left\{\frac{1}{5},\frac{1}{6},\cdots,\frac{1}{10},\frac{1}{12},\cdots,\frac{1}{36},\frac{1}{40},\cdots,\frac{1}{56}\right\}$.
+Figure 3: Results of $\left\|\bm{u}^{k}_{h}\right\|_{H(\mathrm{div})\text{-error}}$, $\left\|P^{k-\frac{1}{2}}_{h}\right\|_{L^{2}\text{-error}}$, $\left\|\bm{\omega}^{k}_{h}\right\|_{H(\mathrm{curl})\text{-error}}$ and $\left\|\bm{H}^{k+\frac{1}{2}}_{h}\right\|_{H(\mathrm{curl})\text{-error}}$ for the spatial convergence tests.
+We use $K\in\left\{12,14,\cdots,32\right\}$ for $N=1$, $K\in\left\{10,12,\cdots,20\right\}$ for $N=2$ and $K\in\left\{8,10,\cdots,14\right\}$ for $N=3$.
+And $\Delta t=\frac{1}{100}$ is employed to avoid pollution of the temporal discretization error, see Fig.
+2.
+Figure 4: Results of $\left\|\nabla\cdot\bm{H}^{k+\frac{1}{2}}_{h}\right\|_{L^{2}}$ for the spatial convergence tests.
+We use $K\in\left\{10,12,\cdots,20\right\}$ for $N=2$, $K\in\left\{8,10,\cdots,14\right\}$ for $N=3$, and $\Delta t=\frac{1}{100}$.
+In this work, the employed finite-dimensional space for $\bm{H}_{h}\in C(\Omega)$ is the mimetic spectral space which is a polynomial space $\mathcal{P}^{N-1,N,N}\times\mathcal{P}^{N,N-1,N}\times\mathcal{P}^{N,N,N-1}$ on orthogonal meshes.
+So we can compute exact divergence of $\bm{H}_{h}$ element-wise, and, when $N=1$, $\nabla\cdot\bm{H}_{h}\equiv 0$.
+For the present results, $\left\|\nabla\cdot\bm{H}^{k+\frac{1}{2}}_{h}\right\|_{L^{2}}:=\sqrt{\sum_{i}H_{i}^{2}}$ where $H_{i}$ is the $L^{2}$-norm of $\nabla\cdot\bm{H}_{h}^{k+\frac{1}{2}}$ in the $i$th element.
+5.1.2 Conservation and dissipation tests
+We test the conservation and dissipation properties of the decoupled formulation in a spatial domain $\Omega=\left(x,y,z\right)\in\left[0,1\right]^{3}$ using manufactured solutions.
+The initial conditions, taken from [5, Section 4.2], are
+$\bm{u}^{0}=\begin{bmatrix}-\sin\left(\pi(x-\frac{1}{2})\right)\cos\left(\pi(y-\frac{1}{2})\right)z\left(z-1\right)\\ \cos\left(\pi(x-\frac{1}{2})\right)\sin\left(\pi(y-\frac{1}{2})\right)z\left(z-1\right)\\ 0\end{bmatrix}$
+and
+$\bm{H}^{0}=\begin{bmatrix}-\sin\left(\pi x\right)\cos\left(\pi y\right)&\cos\left(\pi x\right)\sin\left(\pi y\right)&0\end{bmatrix}^{\mathsf{T}}$
+which possess the initial energy of $\mathcal{E}^{0}\approx 0.25833$ if $\mathsf{c}=1$. The boundary conditions are
+$\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}\quad\text{and}\quad\widehat{P}=0\quad\text{on }\partial\Omega\times(0,T];$
+we use natural boundary conditions on the whole boundary and set them to zero for these tests.
+And the external body force is zero; $\bm{f}=\bm{0}$.
+A mesh of uniform cubic elements and a constant time-step interval is employed.
+The edge length of the element is denoted by $h$.
+The decoupled formulation, (23) and (24), is solved with different $\mathrm{R}_{f}$ and $\mathrm{R}_{m}$.
+The results presented in Fig.
+5 show that mass conservation is always strongly satisfied.
+As for the energy, a dissipation rate of a machine precision error is obtained.
+[…]
+As can be seen in Figure 5, the total energy in the absence of dissipative terms, viscosity, and resistivity is not constant, although it does not deviate too far from its initial value.
+This is due to the fact that $\mathcal{A}_{h}^{k-\frac{1}{2}}-\tilde{\mathcal{A}}^{k-\frac{1}{2}}_{h}\neq 0$ as was shown in Section 4.
+However, the variations in total energy remain bounded, as is expected from a symplectic integrator.
+5.2 Orszag-Tang vortex
+The Orszag-Tang vortex initially studied in [39] is a well-known two-dimensional incompressible MHD test case.
+In a periodic square, the Orszag-Tang vortex gradually develops narrow corridors (that will eventually become singularities) of extreme current density where the magnetic field changes its sign suddenly [40], which makes this test case a challenging one.
+Giving a stream function $\psi$ and a magnetic potential $A$, also see [40, Section 5.3],
+$\psi=2\sin\left(y\right)-2\cos\left(x\right),\quad A=\cos\left(2y\right)-2\cos\left(x\right),$
+the initial conditions are $\bm{u}^{0}=\nabla\times\psi$, and $\bm{H}^{0}=\nabla\times A$.
+The fully orthogonally periodic square is $\Omega=\left(x,y\right)\in\left[0,2\pi\right]^{2}$ and $\bm{f}=\bm{0}$.
+On a mesh of $K^{2}$ uniform cubic elements, we solve the two-dimensional version of the decoupled formulation with a time-step interval $\Delta t=\frac{1}{200}$ and the parameters are $\mathrm{R}_{f}=100$, $\mathrm{R}_{m}=100$, and $\mathsf{c}=1$.
+Some snapshots of $\bm{j}_{h}^{k+\frac{1}{2}}=\nabla\times\bm{H}_{h}^{k+\frac{1}{2}}$ are shown in Fig.
+6, and a good match to reference results in [40] is observed.
+Elementwise $\log_{10}\left(\left|\nabla\cdot\bm{H}_{h}^{k+\frac{1}{2}}\right|\right)$ at $t^{k}=t=1$ for different combinations of $N$ and $K$ are shown in Fig.
+7 where we can see the convergence of weak conservation of Gauss’s law for magnetism under $ph$-refinement.
+Figure 6: $\bm{j}_{h}^{k+\frac{1}{2}}$ at $t^{k}=t\in\left\{0,0.2,0.4,0.6,0.8,1\right\}$ for the Orszag-Tang vortex test using $N=4,K=48,\Delta t=\frac{1}{200}$.
+Figure 7: Element-wise $\log_{10}\left(\left|\nabla\cdot\bm{H}_{h}^{k+\frac{1}{2}}\right|\right)$ at $t^{k}=t=1$ for the Orszag-Tang vortex test using $\Delta t=\frac{1}{200}$.
+Left: $N=2,K=32$.
+Middle: $N=4,K=32$.
+Right: $N=4,K=48$.
+5.3 Lid-driven cavity
+The two-dimensional lid-driven cavity is a domain $\Omega=\left(x,y\right)\in[0,1]^{2}$ with four infinite no-slip solid walls, and there is no external body force; $\bm{f}=\bm{0}$.
+Initially, the fluid is at rest, i.e., $\bm{u}^{0}=\bm{0}$, and is immersed in a magnetic field $\bm{H}^{0}=\begin{bmatrix}0&1\end{bmatrix}^{\mathsf{T}}$.
+At $t>0$, the top wall, namely the lid, moves right with a constant speed $1$, i.e., $\bm{u}_{\mathrm{lid}}=\begin{bmatrix}1&0\end{bmatrix}^{\mathsf{T}}$, and drives the fluid due to viscous entrainment.
+Over the whole boundary, the tangential electric field is zero all the time.
+In summary, we have boundary conditions
+$\bm{u}\cdot\bm{n}=\hat{u}=0,\quad\bm{E}\times\bm{n}=\widehat{E}=0\qquad\text{on }\partial\Omega\times\left(0,T\right],$
+and
+$\bm{u}\times\bm{n}=\hat{u}_{\parallel}=\left\{\begin{aligned} 1\quad&\text{on}\ \Gamma_{y}^{+}\times\left(0,T\right]\\ 0\quad&\text{else}\end{aligned}\right.,$
+where $\Gamma_{y}^{+}$ is the face $\left(x,y\right)\in\left(0,1\right)\times 1$.
+To simulate the lid-driven cavity, an orthogonal spatial mesh of $32^{2}$ elements that are locally refined near the boundary, see Fig.
+8a, is used.
+The time-step interval is $\Delta t=\frac{1}{1000}$, Reynolds numbers are $\mathrm{R}_{f}=\mathrm{R}_{m}=400$, the coupling number is $\mathsf{c}=\mathrm{R}_{m}^{-1}$ (such that $\mathsf{s}=\mathsf{c}\mathrm{R}_{m}=1$) and the polynomial degree is $N=3$.
+The two-dimensional version of the decoupled formulation is solved until a steady state is reached.
+The criterion of the steady state is
+$\frac{1}{\Delta t}\max\left(\left\|\bm{u}_{h}^{k}-\bm{u}_{h}^{k-1}\right\|_{L^{2}},\left\|\bm{H}_{h}^{k+\frac{1}{2}}-\bm{H}_{h}^{k-\frac{1}{2}}\right\|_{L^{2}}\right)<10^{-5}.$
+The flow reaches its steady state at $t=82.049$.
+Some results are presented in Fig.
+8.
+They have a good agreement to the reference results taken from [3].
+In Table 1 and Table 2, solutions along two centerlines are provided as quantitative benchmarks.
+(a) mesh
+(b) $\omega_{h}^{k}$
+(c) streamlines of $\bm{u}_{h}^{k}$
+(d) streamlines of $\bm{H}_{h}^{k+\frac{1}{2}}$
+Figure 8: The mesh and some results at the steady state for the lid-driven cavity test.
+(a): The mesh.
+(b): Vorticity $\omega_{h}^{k}$ for contour lines in $\left\{-5,-4,-3,-2,-1,-0.5,0,0.5,1,2,3,4,5,6,7\right\}$.
+(c): Streamlines of $\bm{u}_{h}^{k}$.
+(d): Streamlines of $\bm{H}_{h}^{k+\frac{1}{2}}$.
+Note that, for diagram (d), a direct numerical integral over a grid of discrete values of $\bm{H}_{h}^{k+\frac{1}{2}}$ is applied.
+Since $\bm{H}_{h}^{k+\frac{1}{2}}$ is not strongly divergence free, there does not exist a stream function $\psi$ such that $\bm{H}_{h}^{k+\frac{1}{2}}=\nabla\times\psi$.
+However, as we have shown that $\bm{H}_{h}^{k+\frac{1}{2}}$ is weakly divergence free for the proposed method, see Fig.
+4 and Fig.
+7, the error in the stream function by this approach is limited.
+Stream functions of contour lines in (c) and (d) are given in See A.
+A good match to reference results in [3, Fig.
+6, Fig.
+7] is obtained.
+Table 1: Solutions along the $x$-direction centerline, i.e., $\left(x,y\right)=[0,1]\times 0.5$, at the steady state for the lid-driven cavity test.
+$\bm{u}_{h}^{k}=\begin{bmatrix}u_{h}^{k}&v_{h}^{k}\end{bmatrix}^{\mathsf{T}}$ and $\bm{H}_{h}^{k+\frac{1}{2}}=\begin{bmatrix}H_{x;h}^{k+\frac{1}{2}}&H_{y;h}^{k+\frac{1}{2}}\end{bmatrix}^{\mathsf{T}}$.
+[…]
+[…]
+0.05 $-0.00152$ $0.11308$ $1.63098$ $0.09510$ $4.90996$
+0.1 $-0.01043$ $0.17267$ $0.63099$ $0.15636$ $4.15657$
+0.15 $-0.02841$ $0.19891$ $0.09245$ $0.14927$ $2.49169$
+0.25 $-0.07561$ $0.22147$ $-0.48799$ $0.13626$ $-0.07002$
+0.5 $-0.14277$ $0.05402$ $-2.19175$ $0.05387$ $-0.07419$
+0.75 $-0.20138$ $-0.25127$ $-2.33726$ $0.41737$ $0.35175$
+0.85 $-0.16772$ $-0.33732$ $0.53077$ $0.73888$ $1.37109$
+0.9 $-0.09707$ $-0.26638$ $2.68539$ $0.56165$ $1.55770$
+0.95 $-0.02697$ $-0.12420$ $3.14441$ $0.27997$ $1.49902$
+1 00 00 $1.72384$ 00 $1.46331$
+Table 2: Solutions along the $y$-direction centerline, i.e., $\left(x,y\right)=0.5\times[0,1]$, at the steady state for the lid-driven cavity test.
+$\bm{u}_{h}^{k}=\begin{bmatrix}u_{h}^{k}&v_{h}^{k}\end{bmatrix}^{\mathsf{T}}$ and $\bm{H}_{h}^{k+\frac{1}{2}}=\begin{bmatrix}H_{x;h}^{k+\frac{1}{2}}&H_{y;h}^{k+\frac{1}{2}}\end{bmatrix}^{\mathsf{T}}$.
+[…]
+[…]
+0.05 $-0.03939$ $-0.00056$ $0.60008$ $-2.40069$ $0.80054$
+0.1 $-0.07135$ $-0.00329$ $0.58025$ $-1.65057$ $0.60078$
+0.15 $-0.11225$ $-0.00676$ $0.70413$ $-0.73480$ $0.39444$
+0.25 $-0.21581$ $-0.00329$ $0.35426$ $0.52230$ $0.06335$
+0.5 $-0.14277$ $0.05402$ $-2.19175$ $0.05387$ $-0.07419$
+0.75 $0.12722$ $0.07170$ $-1.81455$ $-0.11813$ $-0.04774$
+0.85 $0.22351$ $0.05999$ $-1.41003$ $0.33644$ $-0.08662$
+0.9 $0.28772$ $0.04373$ $-2.18569$ $1.71570$ $-0.08702$
+0.95 $0.47275$ $0.01801$ $-6.41153$ $3.05614$ $0.20349$
+1 $1$ 00 $-15.51828$ $-5.29567$ $1$
+6 Conclusion
+In this work, we present a decoupled structure-preserving discretization for the incompressible MHD equations with general boundary conditions.
+A spatially discrete formulation in a mixed finite element setting is first proposed.
+It preserves conservation of mass and conservation of charge strongly, preserves Gauss’s law for magnetism weakly, and preserves the correct energy dissipation rate.
+A leapfrog-type temporal integrator is then applied to the fluid part and the Maxwell part at two staggered time sequences, respectively, such that they are decoupled at the fully discrete level.
+This two-step decoupled method has an optimal spatial accuracy and a second-order temporal accuracy.
+However, the decoupled method does not preserve energy conservation.
+Other invariants, for example, cross- and magnetic-helicity, are preserved by the incompressible MHD equations.
+Possible extensions of this work include the preservation of these quantities.
+Acknowledgments
+The research of Deepesh Toshniwal is supported by project number 212.150 awarded through the Veni research programme by the Dutch Research Council (NWO).
+Andrea Brugnoli acknowledges the financial support of the German Research Foundation (DFG) and the Berlin Mathematics Research Center (MATH+ project AA4-12).
+The research of Yi Zhang is supported by the Natural Science Foundation of Guangxi under grant number 2024JJB110005.
+Appendix A Stream functions
+In Fig. 8c, the streamlines are for stream functions in
+[…]
+[…]
+In Fig. 8d, the streamlines are for stream functions in
+[…]
+[…]
+For both cases, the reference stream function is set to 00 at the bottom-left corner.
+Appendix B A Crank-Nicolson discretization of the coupled formulation
+On the time sequence (22), if we apply the Crank-Nicolson scheme directly to the coupled formulation (11), we will obtain a fully discrete formulation written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, for $k=1,2,3,\cdots$ successively, seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h},\bm{H}^{k}_{h}\right)\in D_{\hat{u}}(\Omega,\Gamma_{\hat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\hat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
+$\left\langle\dfrac{\bm{u}_{h}^{k}-\bm{u}^{k-1}_{h}}{\Delta t},\bm{v}_{h}\right\rangle_{\Omega}+\mathcal{F}\left(\dfrac{\bm{\omega}_{h}^{k}+\bm{\omega}_{h}^{k-1}}{2},\dfrac{\bm{u}_{h}^{k}+\bm{u}_{h}^{k-1}}{2},\dfrac{\bm{H}_{h}^{k}+\bm{H}_{h}^{k-1}}{2},\bm{v}_{h}\right)$
+$-\left\langle{P^{k-\frac{1}{2}}_{h}},{\nabla\cdot\bm{v}_{h}}\right\rangle_{\Omega}$ $=\left\langle{\bm{f}^{k-\frac{1}{2}}},{\bm{v}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}^{k-\frac{1}{2}}},{\mathcal{T}\bm{v}_{h}}\right\rangle_{\Gamma_{\widehat{P}}},$
+$-\left\langle{\bm{u}^{k}_{h}},{\nabla\times\bm{w}_{h}}\right\rangle_{\Omega}+\left\langle{\bm{\omega}^{k}_{h}},{\bm{w}_{h}}\right\rangle_{\Omega}$ $=-\left\langle{\widehat{\bm{u}}^{k}},{\mathcal{T}_{\parallel}\bm{w}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{u}}}},$
+$\left\langle{\nabla\cdot\bm{u}^{k}_{h}},{q_{h}}\right\rangle_{\Omega}$ $=0,$
+$\left\langle{\dfrac{\bm{H}_{h}^{k}-\bm{H}^{k-1}_{h}}{\Delta t}},{\bm{b}_{h}}\right\rangle_{\Omega}+\mathcal{M}\left(\dfrac{\bm{\omega}_{h}^{k}+\bm{\omega}_{h}^{k-1}}{2},\dfrac{\bm{H}_{h}^{k}+\bm{H}_{h}^{k-1}}{2},\bm{b}_{h}\right)$ $=\left\langle{\widehat{\bm{E}}^{k-\frac{1}{2}}},{\mathcal{T}_{\parallel}\bm{b}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}.$
+where $\mathcal{F}$ and $\mathcal{M}$ are functionals
+$\mathcal{F}\left(\bm{\omega}_{h},\bm{u}_{h},\bm{H}_{h},\bm{v}_{h}\right)=a\left(\bm{\omega}_{h},\bm{u}_{h},\bm{v}_{h}\right)+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\bm{\omega}_{h}},{\bm{v}_{h}}\right\rangle_{\Omega}-\mathsf{c}\ a\left(\nabla\times\bm{H}_{h},\bm{H}_{h},\bm{v}_{h}\right),$
+$\mathcal{M}\left(\bm{u}_{h},\bm{H}_{h},\bm{b}_{h}\right)=\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\bm{b}_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\bm{b}_{h}\right).$
+Repeating the analyses in Section 3.2 will reveal that this fully discrete formulation strongly preserves conservation of mass and conservation of charge, and weakly satisfies Gauss’s law for magnetism.
+Especially, compared to the decoupled formulation, it preserves the correct dissipation of energy, which leads to conservation of energy for the ideal MHD (for example, see the bottom-left diagram of Fig.
+5).